Been a while since ive done algebra :)
instructions: Solve each equation. The letter a b and c are constants.
the problem is 1-ax=b and a DOESNT equal 0.

Answer by stanbon(75887)

1-ax=b
-----
1 = ax+b
----
ax = 1-b
----
x = (1-b)a
==============
Cheers,
Stan H.